OpenStudy (anonymous):

Select one of the factors of 5x2 + 7x + 2. (5x - 2) (x + 2) (5x + 1) None of the above @hartnn I want to try and do this one on my own but I was wondering if I get something wrong then you could guide me back?

hartnn (hartnn):

sure, start!

OpenStudy (anonymous):

okay :) lol

OpenStudy (anonymous):

so first you would have to find what makes +2 and 14?

hartnn (hartnn):

\(5x^2 +7x+2\) 2 numbers with SUM = +7 and PRODUCT = 5*2 = 10 ?

OpenStudy (anonymous):

oh picked the wrong number

hartnn (hartnn):

for \(ax^2+bx+c\) it is sum = +b product = ac

OpenStudy (anonymous):

okay lol

hartnn (hartnn):

you can remember as `sum`= middle number product = product of `extreme` numbers

hartnn (hartnn):

** `middle`

OpenStudy (anonymous):

okay i wrote that down :)

OpenStudy (anonymous):

so you would have to find what makes 10 and 7

hartnn (hartnn):

yes sum as 7 and product as 10

OpenStudy (anonymous):

so \[5+2=7 and 5\times2=10\]

hartnn (hartnn):

correct :) so how will we split 7x as ?

OpenStudy (anonymous):

so \[(5+2)(5\times2)?\]

hartnn (hartnn):

we split the middle term right ? we got +5 and +2 so we split 7x as \(\Large 7x = +5x+2x\)

hartnn (hartnn):

\(\Large 5x^2 +5x+2x+2\)

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

so okay now im stuck

hartnn (hartnn):

as always consider first 2 terms \(\Large 5x^2+5x\)

hartnn (hartnn):

what can be factored out ?

OpenStudy (anonymous):

x(5x+5)?

OpenStudy (anonymous):

wait i think i did it wrong

hartnn (hartnn):

x can be factored out, yes but 5 can be factored out too, right ?

OpenStudy (anonymous):

oh yes i was a little confused on that lol so \[5x \left( x+1 \right)\]

hartnn (hartnn):

correct now 2x+2

OpenStudy (anonymous):

2(x+1)

OpenStudy (anonymous):

so it would be............

OpenStudy (anonymous):

\[(5x+2)(x+1)\]

hartnn (hartnn):

excellent! :D

OpenStudy (anonymous):

but what would be the answer tho? im looking at them but idk

OpenStudy (anonymous):

oh wait it would be d

hartnn (hartnn):

its not from the first 3 options so its None :)

hartnn (hartnn):

yes, you got d correct ;)

OpenStudy (anonymous):

thank you for helping me!!!!

hartnn (hartnn):

my pleasure ^_^